OpenStudy (anonymous):

what is -9 = 5/-17N?

OpenStudy (anonymous):

Are we solving for N? If so, you multiply across by -17N so it cancels out on the right side and gets rid of the fraction. This gives us (-9) * (-17N) = 5, when we simplify the left side we get 153N=5. To solve for N we divide by 153 on each side giving us N = 5/153

OpenStudy (anonymous):

Shouldn't it be -153?

OpenStudy (anonymous):

the answer ash.vasvani got is right. remember you are multiplying 2 negatives so that gives a positive

OpenStudy (anonymous):

it will go somewhat like this -9=5/-17N(multiply through by denominator) -9 x -17N=(5/-17N) X -17N 153N=5 (divide by coefficient of N) N=5/153